### v2

@@ Abstract
Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability. Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes. We show that a transformer model can be trained to meet this standard if built using mathematically and logically specified frameworks. In this paper, we fully verify a model for n-digit integer addition. To exhibit the reusability of verified modules, we insert the trained integer addition model into an untrained model and train the combined model to perform both addition and subtraction. We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model. We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of language models built using them. The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety of language models.
@@ 1 Introduction
Figure 1: An overview of our methodology: (1) We trained a perfectly accurate 6-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialisation. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model reuses the inserted addition circuits for both addition and subtraction questions.
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need for understanding their behavior. Many argue that making these models interpretable is key to their safe use (Hendrycks & Mazeika, 2022).
Mechanistic interpretability focuses on demystifying and validating the algorithms behind model weights, translating complex computations into more human-understandable components (Raukur et al., 2022). This understanding aids in predicting model behavior in new situations and fixing model errors.
In creating and training a model, we would ideally want the model to have a very high standard of accuracy and trustworthiness. We aim to achieve this by holding the model to a standard that we term known-good. A model performing a task is known-good if:
@@
    The model’s algorithm for the task and the mechanisms it implements (the “circuits”) are understood.
@@
    All possible edge cases have been identified and tested. As a heuristic, the model should be able to perform the task one million times without error (“1M Q accuracy”).
@@
    The model has no known failure modes.
However, exhaustive testing of a model for most tasks is unfeasible. For instance, a task involving the addition of 5-digit integers (e.g., 12345+67890) presents ten billion possible variations. Rather than testing every variation, the task can be conceptualized within a formal framework to pinpoint edge cases the model must handle. For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases. A known-good model must incorporate algorithms to manage this and other more frequent edge cases. A known-good model would hence have circuits that are verified to perform the task perfectly.
In this paper, we detail the development and interpretation of a known-good model for addition. Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results. We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases. The model achieves a very low training loss and can do one million predictions without error, making it a known-good model.
Additionally, we develop a “mixed” model capable of both addition and subtraction, incorporating the known-good addition model. This mixed model extensively reuses the addition circuits for both operations and learns new circuits, facilitating the interpretation of the model’s algorithm. This approach marks progress toward a known-good model for both addition and subtraction.
@@
  • A trustworthy known-good 5-digit addition model, with an algorithm extensible to n-digits.
  • A 6-digit addition and subtraction model that reuses established addition circuits for both operations, indicating steps toward a known-good model for these tasks.
  • A proof of concept for enhancing the trustworthiness of a larger model by integrating a smaller known-good transformer model.
@@ 2 Related Work
Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020). To better represent these algorithms, Elhage et al. (2021a) introduced a mathematical framework to represent how transformer attention heads can work with each other to implement complex algorithms.
To explain how (not where) a model algorithm is implemented, Jenner et al. (2023) recommends documenting a low-level computation graph (causal abstraction), with a mapping from the graph to the model nodes that implement the computation, with experimentation verification.
Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Stolfo et al. (2023) used causal mediation analysis to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems.
It is also possible to improve a model’s arithmetic abilities through training data enrichment. Liu & Low (2023) used supervised fine tuning (including enriched training data) to convert LLaMA 7B to Goat - a model that is “Good at Arithmetic Tasks” including addition and subtraction. They also showed the model’s performance relied in part on LLaMA’s consistent tokenization of digits.
Studying Toy Models for Arithmetic. Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied. Nanda et al. (2023) studied modular addition in a one-layer transformer (where each number was represented by a distinct token) to better understand training dynamics.
Quirke & Barez (2024) detailed how a 1-layer, 3-head transformer model performed 5-digit addition questions. It showed a rare edge case (e.g. “77778+22222=100000” where a “carry 1” cascades through 4 digits) that the model couldn’t handle - showing the importance of understanding and testing all edge cases for trustworthiness.
Michaud et al. (2023) posits that many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta). Models must learn these quanta to reduce loss. They posit that “understanding a network reduces to enumerating its quanta.” Schaeffer et al. (2023) provides useful ways to measure quanta in mathematical prediction problems.
@@ 3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. Let us define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$. The resulting answer can be represented as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. For an illustrative example, see Figure 2.
@@
Figure 2: We refer to individual tokens in a 5-digit addition question as D4, .. D0, and D’4, .., D’0 and the answer tokens as A5, .., A0. For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
We categorize the sum of each digit pair into three overlapping groups BA, MC and US defined as:
$BA_{n}=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
$MC_{n}=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
$US_{n}=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ 0&\text{otherwise}.\end{cases}$ (3)
When calculating $(A_{n})$, considering only the carry bit from the two preceding digits, we have:
$A_{n}=\begin{cases}D_{n}+D^{\prime}_{n}+1\underbrace{\text{mod}}_{10},&\text{if }\underbrace{MC}_{n-1}=1\text{ or}\\ &(\underbrace{US}_{n-1}=1\text{ and }\underbrace{MC}_{n-2}=1),\\ D_{n}+D^{\prime}_{n}\underbrace{\text{mod}}_{10},&\text{otherwise.}\end{cases}$ (4)
This approach is aligned with the framework implemented by the addition model in Quirke & Barez (2024). However, it is restricted to cases where the carry bit cascades over two or fewer digits. To address this limitation, we introduce an additional digit-level categorization called TriCase:
$D_{n}.C=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\begin{cases}T10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ T9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ T8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8.\end{cases}$ (5)
These categorization cases contain carry bit information:
  • If $D_{n}.C=T10$, the digit-pair always causes a MC,
  • if $D_{n}.C=T9$, the digit-pair will cause an MC if the proceeding digit-pair sum generates an MC, and
  • if $D_{n}.C=T8$, the digit-pair never causes an MC .
To compute the first digit of the answer in a problem such as (55555+44445=)(55555+44445=)( 55555 + 44445 = ), the algorithm must synthesize the five TriCase values from the respective digit-pairs. This synthesis is crucial to determine the presence of a carry bit cascading through all five digits.
We introduce a function, TriAdd, to facilitate the transfer of TriCase data between digits. It is defined as follows:
@@
$\underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C)}=\begin{cases}T10&\text{if }D_{n}.C=T10\\ &\text{or }D_{n}.C=T9\\ &\text{\quad and }D_{n-1}.C=T10,\\ T9&\text{if }D_{n}.C=T8\\ &\text{\quad and }D_{n-1}.C=T10,\\ T8&\text{otherwise.}\end{cases}$ (6)
TriAdd transfers MC data from $(D_{n-1})$ to $(D_{n})$ by integrating the values of $(D_{n}.C)$ and $(D_{n-1}.C)$. This function can be represented as nine bigram mappings and yields three distinct outputs (details in App. D). Finally, we introduce the recursive function $D_{n}$.$C_{m}$ , which aggregates TriCase data across $(m)$ digits to enhance accuracy.
$D_{n}.C_{m}=\begin{cases}\underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C_{m-1})}&\text{if }m>2,\\ \underbrace{\text{TriAdd}}_{(D_{n}.C,D_{n-1}.C)}&\text{if }m=2.\end{cases}$ (7)
Simply put $D_{n}$.$C_{m}$ cascades the carry bit over $m$ digits. Calculating $D_{1}.C_{2}$ means calculating $\text{TriAdd}(D_{1}.C,D_{0}.C)$ combining the $D_{0}$-specific TriCase value and $D_{1}$-specific TriCase value to give a perfectly accurate TriCase value for $D_{1}$. Calculating $D_{4}.C_{5}$ means calculating $\text{TriAdd}(D_{4}.C,D_{3}.C_{4})$ combining the $D_{4}$-specific TriCase value with the perfectly accurate $D_{3}.C_{4}$ TriCase value to give a perfectly accurate TriCase value for $D_{4}$.
In 5-digit addition, the values D4.C5, D3.C4, D2.C3 and D1.C2 are perfectly accurate as they integrate TriCase data all the way back to and including D0.MC . An accurate algorithm must calculate these values. Works on 5-digit model was unable to calculate these $D_{n}$.$C_{m}$ and so was not perfectly accurate (Quirke & Barez, 2024) . As shown later, our model’s algorithm incorporates the (Quirke & Barez, 2024) algorithm, and adds these perfectly accurate $D_{n}$.$C_{m}$ Figure 3 shows the combined algorithm.
Figure 3: In 5-digit addition the model first calculates all $D_{n}.C$ values, then combines them into perfectly accurate $D_{n}$.$C_{m}$ values by P11 (token =) when A5 is predicted. From P12, the remaining answer digits are calculated one by one, using $D_{n}$.$C_{m}$ values if needed. This algorithm can be extended to generalise to n-digits.
3.2 Question Difficulty and Frequency
To analyse question difficulty, we categorised questions by the complexity of the computation required to solve the question, as shown in Tab. 1. The categories are arranged according to the number of digits that a carry bit has to cascade through.
Table 1: We categorise addition questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty (and decreasing occurrence frequency). Five-digit addition questions quanta are S0 to S5. Ten-digit addition question quanta are S0 to S10. S10’s frequency is $\sim 3\times 10^{-4}$ showing the need to enrich training data for rare edge cases.
Name Contains Example Freq
S0 BA 11111+12345=23456 $\sim$5%
S1 BA,MC 11111+9=22230 $\sim$21%
S2 BA,MC,US 11111+89=22300 $\sim$34%
S3 BA,MC,USx2 11111+889=23000 $\sim$28%
S4 BA,MC,USx3 11111+8889=30000 $\sim$11%
S5 BA,MC,USx4 11111+88889=100000 $\sim$2%
@@ 3.3 Techniques
In investigating the circuits of the model, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework. Hence, we define a node as the computation done by an attention head or MLP layer for a given token position.
Mean ablation. To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across its batch.
Attention Patterns. To find out what the model attends to at that node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).

### v7

@@ Abstract
Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability. Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes. We show that a model can be trained to meet this standard if built using mathematically and logically specified frameworks. In this paper, we fully verify an auto-regressive transformer model that performs n-digit integer addition. To exhibit the reusability of verified modules, we insert the trained integer addition model into a larger untrained model and train the combined model to perform both addition and subtraction. We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model. We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of language models built using them. The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of language models.
Increasing Trust in Language Models through the Reuse of Verified Circuits
Philip Quirke Apart Research. philipquirkenz@gmail.com                        Clement Neo Apart Research.                        Fazl Barez Apart Research. Department of Engineering Sciences, University of Oxford fazl@robots.ox.ac.uk
@@ 1 Introduction
Figure 1: An overview of our methodology: (1) We trained an accurate 6-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need for understanding their behavior. Hendrycks and Mazeika (2022) argue that making these models interpretable is key to their safe use. Mechanistic interpretability focuses on demystifying and validating the algorithms behind model weights, translating complex computations into more human-understandable components (Raukur et al., 2022). This understanding aids in predicting model behavior in new situations and fixing model errors.
In creating and training a model, we would ideally want the model to have a very high standard of accuracy and trustworthiness. We aim to achieve this by holding the model to a standard that we term known-good. We define a model performing a task to be known-good if:
@@
    The model’s algorithm for the task and the mechanisms it implements (the “circuits") are understood.
@@
    All possible task edge cases have been identified and tested.
@@
    Empirically, the model prediction accuracy is 99.9999% (a standard reliability measure used in industry and abbreviated as “six nines"). That is, it can perform the task one million times with at most 1 wrong prediction.
Exhaustive testing of a model task may be infeasible. For instance, when adding two 5-digit integers (e.g. 12345+67890) there are ten billion variations. Some tasks can be conceptualized within an existing formal framework that allows identification of all edge cases the model must handle. For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases. A known-good model must incorporate algorithms to manage all known edge cases. A known-good model would hence have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence. circuits that perform the task accurately.
In this paper, we detail the development and interpretation of a known-good model for addition. Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results. We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases. The model achieves a very low training loss and has six nines (99.9999%) accuracy. The model hence achieves our standard of ’known-good’. Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model. This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations and learns new circuits, facilitating the interpretation of the model’s algorithm. This approach marks progress toward a known-good model for both addition and subtraction.
@@
  • Defining several known-good n-digit addition models with six nines accuracy which all use the same algorithm.
  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the more-capable model’s algorithm.
  • Defining several n-digit addition and subtraction models with six nines accuracy, that reuse established addition model circuits for both operations, and detailing progress towards these models being known-good.
@@ 2 Related Work
Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020). Mathematical frameworks (Elhage et al., 2021a) explain how transformer attention heads can work with each other to implement complex algorithms.
Causal Scrubbing (Jenner et al., 2023) recommends explaining a model algorithm by documenting a low-level computation graph, mapping from the graph to the model nodes that implement the computation, and performing experimentation verification. Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems. It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu and Low, 2023).
Studying Toy Models for Arithmetic. Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023). Quirke and Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on rare edge cases (e.g. “77778+22222=100000" where a “carry 1" cascades through 4 digits), highlighting the importance of understanding and testing all edge cases for trustworthiness.
Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized" into discrete chunks (quanta) (Michaud et al., 2023). Models must learn these quanta to reduce loss. Understanding a network reduces to enumerating its quanta. Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
@@ 3 Methodology
A transformer model may learn an addition algorithm that differs significantly from the traditional human process. Here we define an alternative mathematical framework for addition, showing it is equivalent to the traditional process and so is mathematically correct. Then we show that our model implements this alternative framework.
@@ 3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. We define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. Figure 2 shows an illustrative example.
@@
Figure 2: For 5-digit addition, our model has 12 input (question) and 6 output (answer) token positions. We name the question tokens D4, …, D0, and D’4, …, D’0 and the answer tokens A5, …, A0. For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
First, the model will need to sum pairs of digits. As in Quirke and Barez (2024), we define the “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of pairs of question digits, as:
$A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$. The model will also need learn how to calculate whether the addition creates a carry bit. As in Quirke and Barez (2024), we define the "Simple Carry" sub-task $A_{n}.SC$:
$A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
The model in Quirke and Barez (2024) is limited to the framework described so far. While it is capable of handling simple addition and carry bits generated directly from the addition of digit pairs, it encounters difficulties when faced with problems involving ’cascading carry’ bits, wherein a carry bit from one digit position propagates to the next.
To illustrate this limitation, consider the problem "00144+00056=000210". When adding the digits in the tens position, 4+5, no carry bit is generated directly. However, due to the carry bit propagating from the ones position, a carry bit is ultimately required in the tens position. A model that solely calculates the sums of digit pairs and considers carry bits resulting from the direct addition of those pairs will fail to accurately resolve this type of problem. In such cases, it would yield an incorrect result, such as "144+56=110". The 1-layer model could cascade the carry bit up to two digits across, but not more.
3.2 Extending the Mathematical Framework
To answer "44444+55556=" with "100000", an accurate model must predict the first answer digit A5 as "1". To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits. This is especially hard as the model predicts answer tokens from left to right.
We introduce a digit-level sub-task called TriCase that calculates the essential “carry one" information for a single pair of digits $D_{n}$ and $D^{\prime}_{n}$. TriCase has 3 possible outputs representing a definite carry one (ST10), a possible carry one depending on the results of other calculations (ST9), or definitely not a carry one (ST8):
$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8,\\ \end{cases}\end{split}$ (3)
To perform the “cascading carry one" calculation, we introduce a TriAdd sub-task. It handles the case where a possible carry one becomes a definite carry one because the next lower digit pair generated a carry one. TriAdd is defined as:
@@ 3.2 Extending the Mathematical Framework
$\begin{split}A_{n}.SV&=\underbrace{\text{TriAdd}}_{(A_{n}.ST,A_{n-1}.ST)}=\\ &\begin{cases}ST10&\text{if }A_{n}.ST=ST10\text{ or }\\ &\quad(A_{n}.ST=ST9\text{ and }\\ &\quad A_{n-1}.ST=ST10),\\ ST8&\text{otherwise.}\end{cases}\end{split}$ (4)
The model can use $ST$ and $SV$ to accurately calculate A5 as 1 or 0 by using A5 = A4.SV = TriAdd(A4.ST, TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST)))). Note that in calculating A5, the model has also calculated an accurate carry bit for each answer digit. For example, the carry bit for A4 is A3.SV = TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST))).
With this framework, the model only needs the sub-tasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform addition (but some models also use the redundant $A_{n}$.SC sub-task). This framework, if implemented by a model, is sufficient for the model to perform n-digit addition accurately.
Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke and Barez (2024) algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
Figure 3: To predict with 99.9999% accuracy, the addition algorithm first calculates “carry one" values $(A_{n}.ST)$, combining them into “cascading carry one" values $(A_{n}.SV).$ At the “+" token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0. The other answer digits are calculated by combining “base add" $(A_{n}.BA)$ and “carry one" $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
@@ 3.3 Techniques
In investigating the model circuits, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework. Hence, we define a node as the computation done by an attention head or MLP layer for a given token position. To investigate what each node is doing, we use the following techniques:
  1. 1.
    Intervention Ablation. To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across a batch and measure how that impacts downstream performance. We also use (automated, n-digit) intervention ablation tests, specific to each sub-task, to test for the expected sub-task behavior.
  2. 2.
    Attention Patterns. To find out what the model attends to at a node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).
  3. 3.
    Principal Component Analysis (PCA). We use PCA to investigate the outputs of attention heads, especially where our framework suggests the head output may be tri-state or bi-state.
  4. 4.
    Question Complexity. We categorized questions by the complexity of the computation required to solve the question (Refer App. G). The addition categories (named S0 to S4) reflect the number of sequential digits that a “carry one" cascades through. The subtraction categories (named M0 to M4) reflect the number of sequential digits that a “borrow one" cascades through. We analysed which nodes were needed for correct prediction of each category.